# Res-VMamba: Fine-Grained Food Category Visual Classification Using Selective State Space Models with Deep Residual Learning

### v1

@@ 1 Introduction
Food plays a significant role in human life, and with the advent of modern technology and dramatically changed lifestyle and dietary habits, there has been a growing focus on food computingMin et al. (2019a). Within the realm of food computing, food recognition is an important research direction in computer vision and machine learning. In the industry, food image classification can be applied to automating restaurant cooking processes, self-checkout systems, and kitchen waste management. Additionally, food recognition is a vital component in numerous health-related applications, encompassing the nutritional analysis of food and dietary habit management. However, despite ongoing research, there are still several challenges in food image classification.
@@
Fine-grained visual classification represents a formidable task within the field of computer vision, seeking to identify various subcategories within a broader category, such as the various species of birdsWah et al. (2011), Van Horn et al. (2015), medical imagesZhou et al. (2021), carsKrause et al. (2013), Dong et al. (2015), aircraftMaji et al. (2013), dogsDataset (2011), petsParkhi et al. (2012), flowersNilsback and Zisserman (2008), natural imagesHorn et al. (2018), Horn et al. (2021), etc. Food image recognition is also an important branch of Fine-Grained Visual Classification (FGVC)Singla et al. (2016), Fan (2022). The pivotal point in the fine-grained classification is that in addition to learning global features, the model should be able to integrate local features to get global-local information mutually in order to achieve better recognition capabilities. Current methods can be summarized as developing model subnet for localization features or delivering better feature learning methods.
Recently, models based on the State Space Model, such as VMambaLiu et al. (2024a), are regarded as outperforming ViTs on large image datasets like ImageNetDeng et al. (2009). It retains the advantage of capturing both local and global information from input images as ViTs while also enhancing the model speed. However, there is still a lack of research on the application of VMamba to fine-grained datasets. Therefore, this study endeavors to employ VMamba on food images and introduces a model, Res-VMamba, designed specifically for fine-grained datasets. This model incorporates a mechanism to share global and local feature states, aiming to enhance performance on detailed image classification tasks.
A well-defined dataset significantly influences the development of possible research topics and the feature-learning capabilities of models. In this study, we utilized the CNFOOD241 datasetFan (2022), Fan et al. (2023). CNFOOD241 is a Chinese food dataset created by expanding ChineseFoodNetChen et al. (2017), including correcting incorrect labels and increasing the number of images and food categories. In addition to model training, we provided a comparative analysis of CNFOOD241 and other food datasets, illustrating its suitability for research. Unlike other food databases, CNFOOD241 preserves the aspect ratio of images and standardizes the size to 600$\times$600 pixels. This preprocessing step prevents image deformation during data augmentation, which could potentially lead to models learning incorrect semantic features. Furthermore, CNFOOD241 exhibits the relative imbalanced data distribution, making it a more challenging dataset for experimentation. Considering the above factors, we selected CNFOOD241 for experiments on fine-grained food classification.
@@
PFID Chen et al. (2009) 2009 101/4,545 Western $\times$
Food50 Joutou and Yanai (2009a) 2010 50/5,000 Misc. $\times$
Food85 Hoashi et al. (2010) 2010 85/8,500 Misc. $\times$
UEC Food100 Matsuda and Yanai (2012) 2012 100/14,361 Japanese $\surd$
UEC Food256 Kawano and Yanai (2015) 2014 256/25,088 Japanese $\surd$
ETH Food-101 Bossard et al. (2014) 2014 101/101,000 Western $\surd$
Diabetes Anthimopoulos et al. (2014) 2014 11/4,868 Misc. $\times$
UPMC Food-101 Wang et al. (2015) 2015 101/ 90,840 Western $\surd$
Geo-Dish Xu et al. (2015) 2015 701/117,504 Misc. $\surd$
UNICT-FD889 Farinella et al. (2015) 2015 889/3,583 Misc. $\surd$
Vireo Food-172 Chen and Ngo (2016) 2016 172/110,241 Chinese $\surd$
Food-975 Zhou and Lin (2016) 2016 975/37,785 Misc. $\times$
Food500 Merler et al. (2016) 2016 508/148,408 Misc. $\times$
Food11 Singla et al. (2016) 2016 11/16,643 Misc. $\surd$
UNICT-FD1200 Farinella et al. (2016) 2016 1,200/4,754 Misc. $\surd$
Food524DB Ciocca et al. (2017) 2017 524/247,636 Misc. $\times$
ChineseFoodNet Chen et al. (2017) 2017 208/192,000 Chinese $\surd$
Vegfru Hou et al. (2017) 2017 292/160,000 Misc. $\surd$
Sushi-50 Qiu et al. (2019) 2019 50/3,963 Japanese $\surd$
FoodX-251 Kaur et al. (2019) 2019 251/158,846 Misc. $\surd$
ISIA Food-200 Min et al. (2019b) 2019 200/197,323 Misc. $\surd$
FoodAI-756 Sahoo et al. (2019) 2019 756/400,000 Misc. $\times$
Taiwanese-Food-101 Yang (2020) 2020 101/20200 Chinese $\surd$
ISIA Food-500 Min et al. (2020) 2020 500/399,726 Misc. $\surd$
Food2K Min et al. (2023) 2021 2,000/1,036,564 Misc. $\surd$
MyFoodRepo-273 SP et al. (2022) 2022 273/24,119 Misc. $\times$
CNFOOD-241 Fan et al. (2023) 2022 241/191,811 Chinese $\surd$
@@ 2.1 Food Recognition Datasets
In the burgeoning field of food computation, the proliferation of food datasets has marked a significant advancement, drawing widespread academic and practical interest. From the inception of datasets like ETH Food-101Bossard et al. (2014), which introduced over one hundred thousand images of Western food varieties, to the expansive collections of ISIA Food-500Min et al. (2020) and Food2KMin et al. (2023), encompassing nearly four hundred thousand and over a million images respectively, the evolution is notable. These datasets, predominantly sourced through web scraping, have been instrumental in advancing computational gastronomy and nutrition studies. However, they share a critical limitation: a lack of uniformity in the size distribution of images across different categories. This variance can lead to substantial discrepancies in some categories, where a few images might significantly exceed the average size of others, potentially skewing the dataset’s overall utility and introducing biases in the processing and classification results obtained after resizing images for analysis.
@@
In response to these challenges, our search for a more consistent and high-resolution dataset led us to CNFOOD-241. Among publicly available food datasets with uniform image sizes, such as UNICT-FD889Farinella et al. (2015), Vireo Food-172Chen and Ngo (2016), and UNICT-FD1200Farinella et al. (2016), CNFOOD-241 distinguishes itself by offering the highest resolution. This characteristic renders it an exceptional resource for conducting detailed image analyses within the food computation domain, facilitating more accurate and reliable studies in food recognition, nutritional analysis, and other related areas. The unveiling of the CNFOOD-241 dataset marks a significant advancement in fulfilling the essential demand for high-quality, uniform datasets within the domain of food computation, thereby facilitating novel pathways for research and innovation. Experimental evidence further indicates that this dataset presents a considerable challenge, enhancing its benchmark value and making it a critical asset for advancing the state of the art in food recognition technologies. This complexity not only tests the limits of current models but also underscores the dataset’s potential as a valuable tool for benchmarking advancements in the field.
@@ 2.2 Fine-Grained Visual Classification
In the domain of computer vision, food recognition is categorically placed within the realm of FGVC, a field distinguished by its focus on distinguishing between closely related subcategories within a broader category. This area has seen the development and application of several state-of-the-art (SOTA) models, each contributing to advancements in dataset-specific performance. Among these, the CMAL-NetLiu et al. (2023) stands out, having been constructed by integrating three expert modules with a CNN-based backbone. Each expert module processes feature maps from specific layers, delivering both a categorical prediction and an attention region. This attention region not only highlights areas of interest within the images but also serves as a means for data augmentation for the other expert modules, thereby enhancing the model’s overall accuracy and robustness.
Furthermore, EfficientNetMingxing Tan (2019), a model that has retained its relevance over time, employs Network Architecture Search (NAS) to devise a set of rules that are consolidated into Compound Model Scaling. This approach strategically scales model architecture to optimize performance. Additionally, ConvNeXTLiu et al. (2022), inspired by the Swin TransformerLiu et al. (2021) architecture, reimagines CNNs to surpass the Swin Transformer’s performance on ImageNet, marking a significant achievement in model design.
Another noteworthy development is the introduction of the Squeeze-and-Excitation (SE) block into ViT, leading to the creation of RepViTWang et al. (2023). This model emerges as a successor in the ViT era following the SENetHu et al. (2018), demonstrating impressive performance in image recognition tasks. Traditional ViTs, which marked a pivotal moment by outperforming models with purely CNN-based backbones in image recognition domains, further underscore the ongoing evolution and innovation within FGVC. These advancements not only push the boundaries of what’s achievable in fine-grained classification but also provide a rich tapestry of methodologies for addressing the nuanced challenges of food recognition.
@@ 2.3 Food Image Recognition
Early food recognition systems primarily used traditional machine learning algorithms. Researchers extract handcrafted features from images using methods such as color histogramsJoutou and Yanai (2009b), Scale-Invariant Feature Transform (SIFT), Histogram of Oriented Gradients (HOG)Ravì et al. (2015), Gabor texturesJoutou and Yanai (2009b), and Local Binary Pattern (LBP)Ravì et al. (2015). These extracted features were then fed into classifiers such as SVMHearst et al. (1998) for categorization. While achieving reasonable performance, these early methods relied heavily on manual feature engineering and were limited by image quality and variability in food appearances.
The emergence of deep learning revolutionized food identification research. Researchers began applying convolutional neural networks, such as AlexNetRahmat and Kutty (2021), ResNet-50Zahisham et al. (2020), EfficientNetG. et al. (2022), and Inception v3Hassannejad et al. (2016), to food image data using transfer learning methods. This approach eliminated the need for manual feature extraction and allowed models to learn hierarchical visual representations from large-scale labeled datasets. Subsequent research enhanced food recognition capabilities using ensemble networks, multi-task learning, and other techniques. Notably, PRENetMin et al. (2023) emerged as a milestone in food recognition, integrating three different branches, each tailored for capturing different aspects of food images. By fusing features from low and high-level layers, PRENet achieved SOTA performance on CNN models.
@@ 2.4 State Space Model on Visual Recognition
Recent research predominantly utilizes CNNs and ViTs for the task of classifying food categories. However, the capability of these models to detect features has not reached the SOTA in large-scale image classification challenges lately, being outperformed by a new generation of models known as Structured State Space for Sequence (S4)Gu et al. (2022) modeling. The improvement of S4 models with a Selection mechanism and their execution using a Scan (S6)Albert Gu (2023), informally known as Mamba, has shown to outclass the Transformer architecture in handling long sequences. The VMambaLiu et al. (2024a) model, which applies the Mamba mechanism to downstream image tasks like classification, now sets the SOTA benchmark on the ImageNet dataset. There are several Mamba models used on vision related task such as VMamba, Vision MambaZhu et al. (2024) have tried to use Mamba to do visual downstream tasks like image classification and object detection, but Vision Mamba more focus on inference speed and GPU-memory usage efficiency. U-MambaJun Ma (2024), VM-UNetJiacheng Ruan (2024), Swin-UMambaLiu et al. (2024b), nnMambaGong et al. (2024) and MambaMorphTao Guo (2024) have replaced convolutional blocks or downsample blocks to Mamba blocks applied on medical image segmentation tasks. VivimYijun Yang (2024) used Mamba-based backbone for medical video object segmentation. However, there is still lack of utiling VMamba on eithor fine-grained data or food recognition task, in this work we propose the VMamba-based model into these downstream task.
@@ 2.5 Deep Residual Learning on Space State Model
Deep residual learning is a technique employed in training deep neural networks that offers several notable advantages introduced by ResNetHe et al. (2016). Its primary advantage lies in addressing challenges such as vanishing gradients and exploding gradients, which commonly impede the training of deep networks. By introducing residual blocks and skip connections, deep residual learning facilitates the flow of gradients throughout the network, effectively mitigating the issue of vanishing gradients. Consequently, it enables the training of deeper neural networks without sacrificing performance. Deeper networks afford the extraction of more complex features, thereby enhancing the model’s representational capacity. Additionally, the training process in deep residual learning converges more efficiently due to expedited gradient propagation via skip connections, resulting in reduced training time and computational costs. Moreover, the technique enhances the model’s generalization capabilities by effectively fitting training data while mitigating the risk of overfitting.
@@ 3.2 Discretization
The primary aim of discretization is to transmute the continuous ODE into a discrete function. This conversion is vital for aligning the model with the input data’s sample rate, thereby enabling computationally efficient operations Gu et al. (2021). Given the input $x_{k}\in\mathbb{R}^{L\times D}$, which is a sampled vector from the signal sequence of length $L$, the ODE Gupta et al. (2022) (Eq.1) can be discretized employing the zeroth-order hold approach:
@@
where $A_{d}=e^{A\Delta}$, $B_{d}=(e^{A\Delta}-I)A^{-1}B$, and $C_{d}=C$, with $B,C\in\mathbb{R}^{D\times N}$ and $\Delta\in\mathbb{R}^{D}$. Following the practice in Albert Gu (2023), the approximation of $B$ through first-order Taylor series is refined as:
@@ 3.4 VMamba Model
The overall architecture of VMamba Model is illustrated in Liu et al. (2024a). The VMamba architecture, showcased in Figure 1, commences by partitioning the input image into patches through a stem module, emulating ViTs while maintaining the 2D structure and translating the patches into a 1D sequence. This approach yields a feature map with dimensions $\frac{H}{4}\times\frac{W}{4}\times C_{1}$.
VMamba layers a series of VSS blocks like Figure 2 atop this feature map to construct “Stage 1,” preserving its dimensions. Hierarchical structures in VMamba are established via down-sampling in “Stage 1” through a patch merging process, as described in Liu et al. (2024a). More VSS blocks are then integrated, reducing the output resolution to $\frac{H}{8}\times\frac{W}{8}$ for “Stage 2.” This down-sampling is reiterated to form “Stage 3” and “Stage 4,” with resolutions of $\frac{H}{16}\times\frac{W}{16}$ and $\frac{H}{32}\times\frac{W}{32}$, respectively. The resulting hierarchical design mirrors the multi-scale representation characteristic of renowned CNN models and some ViTs. Thus, VMamba’s architecture emerges as a comprehensive and versatile candidate for a variety of vision-related applications with analogous requirements.

### v2

@@ 1 Introduction
Food plays a significant role in human life, and with the advent of modern technology and dramatically changed lifestyle and dietary habits, there has been a growing focus on food computing [1]. Within the realm of food computing, food recognition is an important research direction in computer vision and machine learning. In the industry, food image classification can be applied to automating restaurant cooking processes, self-checkout systems, and kitchen waste management. Additionally, food recognition is a vital component in numerous health-related applications, encompassing the nutritional analysis of food and dietary habit management. However, despite ongoing research, there are still several challenges in food image classification.
@@
Fine-grained visual classification represents a formidable task within the field of computer vision, seeking to identify various subcategories within a broader category, such as the various species of birds [2], [3], medical images [4], cars [5], [6], aircraft [7], dogs [8], pets [9], flowers [10], natural images [11], [12], etc. Food image recognition is also an important branch of Fine-Grained Visual Classification (FGVC) [13], [14]. The pivotal point in the fine-grained classification is that in addition to learning global features, the model should be able to integrate local features to get global-local information mutually in order to achieve better recognition capabilities. Current methods can be summarized as developing model subnet for localization features or delivering better feature learning methods.
Recently, models based on the State Space Model, such as VMamba [15], are regarded as outperforming ViTs on large image datasets like ImageNet [16]. It retains the advantage of capturing both local and global information from input images as ViTs while also enhancing the model speed. However, there is still a lack of research on the application of VMamba to fine-grained datasets. Therefore, this study endeavors to employ VMamba on food images and introduces a model, Res-VMamba, designed specifically for fine-grained datasets. This model incorporates a mechanism to share global and local feature states, aiming to enhance performance on detailed image classification tasks.
A well-defined dataset significantly influences the development of possible research topics and the feature-learning capabilities of models. In this study, we utilized the CNFOOD241 dataset [14], [17]. CNFOOD241 is a Chinese food dataset created by expanding ChineseFoodNet [18], including correcting incorrect labels and increasing the number of images and food categories. In addition to model training, we provided a comparative analysis of CNFOOD241 and other food datasets, illustrating its suitability for research. Unlike other food databases, CNFOOD241 preserves the aspect ratio of images and standardizes the size to 600$\times$600 pixels. This preprocessing step prevents image deformation during data augmentation, which could potentially lead to models learning incorrect semantic features. Furthermore, CNFOOD241 exhibits the relative imbalanced data distribution, making it a more challenging dataset for experimentation. Considering the above factors, we selected CNFOOD241 for experiments on fine-grained food classification.
@@
PFID [19] 2009 101/4,545 Western $\times$
Food50 [20] 2010 50/5,000 Misc. $\times$
Food85 [21] 2010 85/8,500 Misc. $\times$
UEC Food100 [22] 2012 100/14,361 Japanese $\surd$
UEC Food256 [23] 2014 256/25,088 Japanese $\surd$
ETH Food-101 [24] 2014 101/101,000 Western $\surd$
Diabetes [25] 2014 11/4,868 Misc. $\times$
UPMC Food-101 [26] 2015 101/ 90,840 Western $\surd$
Geo-Dish [27] 2015 701/117,504 Misc. $\surd$
UNICT-FD889 [28] 2015 889/3,583 Misc. $\surd$
Vireo Food-172 [29] 2016 172/110,241 Chinese $\surd$
Food-975 [30] 2016 975/37,785 Misc. $\times$
Food500 [31] 2016 508/148,408 Misc. $\times$
Food11 [13] 2016 11/16,643 Misc. $\surd$
UNICT-FD1200 [32] 2016 1,200/4,754 Misc. $\surd$
Food524DB [33] 2017 524/247,636 Misc. $\times$
ChineseFoodNet [18] 2017 208/192,000 Chinese $\surd$
Vegfru [34] 2017 292/160,000 Misc. $\surd$
Sushi-50 [35] 2019 50/3,963 Japanese $\surd$
FoodX-251 [36] 2019 251/158,846 Misc. $\surd$
ISIA Food-200 [37] 2019 200/197,323 Misc. $\surd$
FoodAI-756 [38] 2019 756/400,000 Misc. $\times$
Taiwanese-Food-101 [39] 2020 101/20200 Chinese $\surd$
ISIA Food-500 [40] 2020 500/399,726 Misc. $\surd$
Food2K [41] 2021 2,000/1,036,564 Misc. $\surd$
MyFoodRepo-273 [42] 2022 273/24,119 Misc. $\times$
CNFOOD-241 [17] 2022 241/191,811 Chinese $\surd$
@@ 2.1 Food Recognition Datasets
In the burgeoning field of food computation, the proliferation of food datasets has marked a significant advancement, drawing widespread academic and practical interest. From the inception of datasets like ETH Food-101 [24], which introduced over one hundred thousand images of Western food varieties, to the expansive collections of ISIA Food-500 [40] and Food2K [41], encompassing nearly four hundred thousand and over a million images respectively, the evolution is notable. These datasets, predominantly sourced through web scraping, have been instrumental in advancing computational gastronomy and nutrition studies. However, they share a critical limitation: a lack of uniformity in the size distribution of images across different categories. This variance can lead to substantial discrepancies in some categories, where a few images might significantly exceed the average size of others, potentially skewing the dataset’s overall utility and introducing biases in the processing and classification results obtained after resizing images for analysis.
@@
In response to these challenges, our search for a more consistent and high-resolution dataset led us to CNFOOD-241. Among publicly available food datasets with uniform image sizes, such as UNICT-FD889 [28], Vireo Food-172 [29], and UNICT-FD1200 [32], CNFOOD-241 distinguishes itself by offering the highest resolution. This characteristic renders it an exceptional resource for conducting detailed image analyses within the food computation domain, facilitating more accurate and reliable studies in food recognition, nutritional analysis, and other related areas. The unveiling of the CNFOOD-241 dataset marks a significant advancement in fulfilling the essential demand for high-quality, uniform datasets within the domain of food computation, thereby facilitating novel pathways for research and innovation. Experimental evidence further indicates that this dataset presents a considerable challenge, enhancing its benchmark value and making it a critical asset for advancing the state of the art in food recognition technologies. This complexity not only tests the limits of current models but also underscores the dataset’s potential as a valuable tool for benchmarking advancements in the field.
@@ 2.2 Fine-Grained Visual Classification
In the domain of computer vision, food recognition is categorically placed within the realm of FGVC, a field distinguished by its focus on distinguishing between closely related subcategories within a broader category. This area has seen the development and application of several state-of-the-art (SOTA) models, each contributing to advancements in dataset-specific performance. Among these, the CMAL-Net [43] stands out, having been constructed by integrating three expert modules with a CNN-based backbone. Each expert module processes feature maps from specific layers, delivering both a categorical prediction and an attention region. This attention region not only highlights areas of interest within the images but also serves as a means for data augmentation for the other expert modules, thereby enhancing the model’s overall accuracy and robustness.
Furthermore, EfficientNet [44], a model that has retained its relevance over time, employs Network Architecture Search (NAS) to devise a set of rules that are consolidated into Compound Model Scaling. This approach strategically scales model architecture to optimize performance. Additionally, ConvNeXT [45], inspired by the Swin Transformer [46] architecture, reimagines CNNs to surpass the Swin Transformer’s performance on ImageNet, marking a significant achievement in model design.
Another noteworthy development is the introduction of the Squeeze-and-Excitation (SE) block into ViT, leading to the creation of RepViT [47]. This model emerges as a successor in the ViT era following the SENet [48], demonstrating impressive performance in image recognition tasks. Traditional ViTs, which marked a pivotal moment by outperforming models with purely CNN-based backbones in image recognition domains, further underscore the ongoing evolution and innovation within FGVC. These advancements not only push the boundaries of what’s achievable in fine-grained classification but also provide a rich tapestry of methodologies for addressing the nuanced challenges of food recognition.
@@ 2.3 Food Image Recognition
Early food recognition systems primarily used traditional machine learning algorithms. Researchers extract handcrafted features from images using methods such as color histograms [49], Scale-Invariant Feature Transform (SIFT), Histogram of Oriented Gradients (HOG) [50], Gabor textures [49], and Local Binary Pattern (LBP) [50]. These extracted features were then fed into classifiers such as SVM [51] for categorization. While achieving reasonable performance, these early methods relied heavily on manual feature engineering and were limited by image quality and variability in food appearances.
The emergence of deep learning revolutionized food identification research. Researchers began applying convolutional neural networks, such as AlexNet [52], ResNet-50 [53], EfficientNet [54], and Inception v3 [55], to food image data using transfer learning methods. This approach eliminated the need for manual feature extraction and allowed models to learn hierarchical visual representations from large-scale labeled datasets. Subsequent research enhanced food recognition capabilities using ensemble networks, multi-task learning, and other techniques. Notably, PRENet [41] emerged as a milestone in food recognition, integrating three different branches, each tailored for capturing different aspects of food images. By fusing features from low and high-level layers, PRENet achieved SOTA performance on CNN models.
@@ 2.4 State Space Model on Visual Recognition
Recent research predominantly utilizes CNNs and ViTs for the task of classifying food categories. However, the capability of these models to detect features has not reached the SOTA in large-scale image classification challenges lately, being outperformed by a new generation of models known as Structured State Space for Sequence (S4) [56] modeling. The improvement of S4 models with a Selection mechanism and their execution using a Scan (S6) [57], informally known as Mamba, has shown to outclass the Transformer architecture in handling long sequences. The VMamba [15] model, which applies the Mamba mechanism to downstream image tasks like classification, now sets the SOTA benchmark on the ImageNet dataset. There are several Mamba models used on vision related task such as VMamba, Vision Mamba [58] have tried to use Mamba to do visual downstream tasks like image classification and object detection, but Vision Mamba more focus on inference speed and GPU-memory usage efficiency. U-Mamba [59], VM-UNet [60], Swin-UMamba [61], nnMamba [62] and MambaMorph [63] have replaced convolutional blocks or downsample blocks to Mamba blocks applied on medical image segmentation tasks. Vivim [64] used Mamba-based backbone for medical video object segmentation. However, there is still lack of utiling VMamba on eithor fine-grained data or food recognition task, in this work we propose the VMamba-based model into these downstream task.
@@ 2.5 Deep Residual Learning on Space State Model
Deep residual learning is a technique employed in training deep neural networks that offers several notable advantages introduced by ResNet [65]. Its primary advantage lies in addressing challenges such as vanishing gradients and exploding gradients, which commonly impede the training of deep networks. By introducing residual blocks and skip connections, deep residual learning facilitates the flow of gradients throughout the network, effectively mitigating the issue of vanishing gradients. Consequently, it enables the training of deeper neural networks without sacrificing performance. Deeper networks afford the extraction of more complex features, thereby enhancing the model’s representational capacity. Additionally, the training process in deep residual learning converges more efficiently due to expedited gradient propagation via skip connections, resulting in reduced training time and computational costs. Moreover, the technique enhances the model’s generalization capabilities by effectively fitting training data while mitigating the risk of overfitting.
@@ 3.2 Discretization
The primary aim of discretization is to transmute the continuous ODE into a discrete function. This conversion is vital for aligning the model with the input data’s sample rate, thereby enabling computationally efficient operations [66]. Given the input $x_{k}\in\mathbb{R}^{L\times D}$, which is a sampled vector from the signal sequence of length $L$, the ODE [67] (Eq.1) can be discretized employing the zeroth-order hold approach:
@@
where $A_{d}=e^{A\Delta}$, $B_{d}=(e^{A\Delta}-I)A^{-1}B$, and $C_{d}=C$, with $B,C\in\mathbb{R}^{D\times N}$ and $\Delta\in\mathbb{R}^{D}$. Following the practice in [57], the approximation of $B$ through first-order Taylor series is refined as:
@@ 3.4 VMamba Model
The overall architecture of VMamba Model is illustrated in [15]. The VMamba architecture, showcased in Figure 1, commences by partitioning the input image into patches through a stem module, emulating ViTs while maintaining the 2D structure and translating the patches into a 1D sequence. This approach yields a feature map with dimensions $\frac{H}{4}\times\frac{W}{4}\times C_{1}$.
VMamba layers a series of VSS blocks like Figure 2 atop this feature map to construct “Stage 1,” preserving its dimensions. Hierarchical structures in VMamba are established via down-sampling in “Stage 1” through a patch merging process, as described in [15]. More VSS blocks are then integrated, reducing the output resolution to $\frac{H}{8}\times\frac{W}{8}$ for “Stage 2.” This down-sampling is reiterated to form “Stage 3” and “Stage 4,” with resolutions of $\frac{H}{16}\times\frac{W}{16}$ and $\frac{H}{32}\times\frac{W}{32}$, respectively. The resulting hierarchical design mirrors the multi-scale representation characteristic of renowned CNN models and some ViTs. Thus, VMamba’s architecture emerges as a comprehensive and versatile candidate for a variety of vision-related applications with analogous requirements.